OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (4, 5) and (6, 8) A. y equals two-thirds x plus seven-thirds B. y equals three-halves x plus 11 C. y equals three-halves x minus 1 D. y equals three-halves minus 8

OpenStudy (anonymous):

It would be C use the slope intercept formula y=mx+b

OpenStudy (anonymous):

The equation of a line is (y-y1) = m(x-x1) You can find m with the slope formula m = (y2-y1)/(x2-x1) = rise / run Let's find m first. m = (8-5)/(6-4) = 3/2 Next plug in the values to the equation of a line y - 5 = 3/2 (x - 4) y = 3/2 x - 6 + 5 y = 3/2 x - 1